OpenStudy (anonymous):

Slove for y ? y^2-6y+5=0

OpenStudy (anonymous):

y² - 6y + 5 = 0 Factor. (y - 5)(y - 1) = 0 Zero product property. y = 1, 5 If you don't know what the zero product property is: Anything times 0 is 0. Right? Well, can we get to zero any other way? Nope, at least one of the numbers we're multiply must be zero. So, if I tell you that ab = 0, you can say either a = 0 or b = 0. That's what the zero product property is. Either y - 5 = 0 or y - 1 = 0. You just need to solve those two simple equations to find what y can be.